Reports 1-1 of 1 Clear search Modify search
LAS (General)
shinji.miyoki - 15:29 Saturday 16 August 2025 (34840) Print this report
FB laser seems to be out of order

Abstract

The FB laser seems to be out of order. we will discuss the recovery plan.

 

Contents

Around 5AM, Ushiba-kun noticed that PMC could not be locked from the last night for over 6 hours, and no flashing in the PMC trans.  At this moment,

  • FB laser output  of 19W was displayed. PMC REFL image of the GigE camera is OK. The transmitting light of the PMC could be faintly visible.
  • PSL room temperature had no changes.
  • The last unlock had no correlation to seismic noise enhancement cause by earthquakes.

So, firstly we suspected that the lock loss happened because of the control side events, such as some troubles in the PMC control including PZT itself.

We discussed what we should do among Ushiba-kun, Miyakawa-kun, Ikeda-san and Miyoki around 9AM via zoom1, and Miyakawa-kun decided to enter the PSL room to check thes present situation.  The followings are waht Miyakawa-kun checked.

  • Displayed values of the HV amp in the PSL room were OK.
  • Turned off the HV amp, and disconnected the cable to the PZT at the output of the HV amp, then
    • Check the resistance of PZT at the terminals on the PMC body. Over Mega-ohm was measured. ---> No problem.
      • When he measured the resisstance of the PZT with a multi-meter, he observed the resistance increased up to over Mega-ohm for ~2 seconds. This means that the capacitance did exist at the measuring point. So coplete disconnection to the PZT path did not happen.
      • Shorting between the PZT cables with a wire, then measured the resistance again. The resistance became over Mega-ohm again for ~ 5 seconds that is longer than the initial ~ 2 seconds.  
      • Touch the cables from the terminal to the PZT by finger, then measured the PZT resistance again. Over Mega-ohm was also measured.
  • HV check
    • Turned on the HV amp and check the voltage at its output. 150V was measured ---. No problem.
  • Cable check
    • Measured the resistance between the signal line and return line of the cable between the HV amp output and the terminal on the PMC, then no shoting was conffirmed. 
  • Reconnect the cables from the HV amp to the PZT again.
    • Check voltage at the terminal on the PMC, then 150V was measured ---> No problem.
  • Beam paths
    • PMC trance beam was confirmed, not only just after the PMC but also the IMC REFL.
    • A beam at the back of the mirror (named C) just beside the input mirror (named A) was confirmed.
    • A beam at the back of the mirror (named D)just beside the output mirror (named B) was confirmed.
    • So, misalignment of the A and D mirrors were possible. But no way to checck it. 
  • PZT scan again
    • No flash was detected.

At this point we could not find any reason why the PMC cannot flash by the PZT scan.

Finally, Miyakawa-kun found the strange situation in the FB laser as,

  • The seed laser was off (red indicator in the front pannel). However, the FB amp was on. In the control application in the PC showed, 19W at the FB output and 1.8W seed power.

However, this situation is the worst case that definetely would result in disorder of the FB laser. So, Miyakawa-kun reduced the current for the FB amp, then actually, the output power of the FB laser reduced and the GigE camera image of the PMC refl became smaller, then finally vanished when the current became zero.

After that Miyakawa-kun tried to recover the master laser, then it came back. At this point, the master laser power was dected in the controller application to be 1.5W. After that he increased the current of the FB amp. However, no output from the FBL. We tried the full process for the FB amp recovery, incuding the windos PC reboot. However, FBL output could not be recovered.

Finally, we suspected the disorder of the FBL.

 

Consideration

  • Why the IR of 19W was emitted without no seed power?
    • We expected the beam of the FBL did not come out when the seed light was off. however, it was emitted.
  • Why the PMC flash vanished even if 19W laser was injected in to the PMC
    • We suspected that the FBL became no-coherence light and multi-mode oscillation happened. So no way to resonate in the PMC?
Comments to this report:
osamu.miyakawa - 15:29 Saturday 16 August 2025 (34841) Print this report

Some detailed process for checking the laser:

When we tried to change the wavelength by changing the temperature of the laser crystal, I noticed that the master laser was off. Even so, the output of the laser power amplifier was still around 19W. This was quite strange, as normally, when the master is off, the power amplifier should also be off by the software interlock system. In addition, this situation, where the master is off but the power amplifier is on, is a situation that could cause damage to the amplifier. When I checked the photo of the control software taken at that time, even though the master was off, the software was showing a value of about 1.6, which is the same as when the master is on. It seems that the amplifier was not recognizing that the master had failed. Note that when I looked at the amplifier output with a sensor card, I could not see any light on the card, so it has been confirmed that the master was not actually producing any light.

First, I reduced the amplifier output to 0 and switched off the power by turning the key on the controller. To be sure, I also switched off the power by turning the key of the master controller. The control software was still running, so it seemed that the connection was still active. I turned the master controller key, turned the amplifier key, pressed the green button on the master controller to turn it on, and started oscillation. When I checked the amplifier output with the sensor card, I could see a faint light, so I thought the master was producing some light with the indicator of ~0.9A. Then, I tried to enable the power amplifier on the software, but it warned that “Mater faulted”, and wouldn’t go further. Once again, I turned off the master and amplifier, then unplugged the power cable of the amplifier, and shut down the software. After that, I turned the master and amplifier keys, launched the software, connected the software to the amplifier, turned on the master by pressing the green button, and pressed the "Enable" button on the software controller. And this time the display showed "Laser ON". 

However, even when I increased the amplifier's current from this point, there was no change in the power at the amplifier's output monitor. However, even when I increase the amplifier's current from this point, there is no change in the power of the amplifier's output monitor. Normally, it should show around 1W when enabled, but it stayed at 0W, and even when I increased the current to around 6A, it was still at 0W.

At this point, I restarted all the devices, then restarted the Windows computer that was running the amplifier control software, and repeated the same procedure, but the result was the same. Even when I increased the amplifier's current, the actual power did not increase. Just to be sure, I checked the output side with a sensor card, but the brightness was the same as when only the master was turned on. So we concluded that the amplifier was not working.

After about 30 minutes, just to be sure, I restarted all the laser devices one more time, but the result was the same. Unless there is something else I am forgetting, it seems likely that the power amplifier is broken.

satoru.ikeda - 15:29 Saturday 16 August 2025 (34842) Print this report

9:20 Shutters open 
9:24 Entry
9:26 Parking
9:27 Enter center
9:34 Entering PSL
 9:45 PZT power off
 9:46 Unplug cable, measure resistance
 9:47 Hi vol output on, measure voltage
 9:54 Measure cable resistance
 9:59 Connect cable, power ON
 10:23 Hivol power off, disconnect PZT
 10:24 Short between PZT terminals with cable
 10:26 Check GND with optical surface plate for short
 10:31 Check resistance value changes by touching cable
 10:45 Cable connected, power on, 150v
 10:47 PMC temperature adjustment value clear
 == Laser ==
 10:55 Reduce laser power, stop amplifier
 11:00 Turn master on
 11:02 Master enable, master off
 11:03 Check interlock
 11:04 Power amp off, master off
 11:05 Turn master key
 11:06 Then turn amp too
 11:06 Turn master on
 11:07 Master enable, output Turn up
 11:08 Turn master off, amp off too
 11:09 Close software, unplug power cable too
 11:10 Amp power off
 11:11 Start software, connect, turn master on
 11:12 Amp on, second stage on, seed?on
 11:14 Master enable
 11:30 Turn off again and restart
 11:39 Turn off master, amps, etc. and disconnect power
11:41 Leave PSL
11:50 Leave center
11:51 Leave parking
11:54 Shutters closed 

shinji.miyoki - 15:36 Saturday 16 August 2025 (34843) Print this report

Mio-sensei pointed out the disorder of the master laser itself.

shinji.miyoki - 17:21 Saturday 16 August 2025 (34845) Print this report

Sorry for my mistake as follows,

shinji.miyoki - 18:22 Saturday 16 August 2025 (34847) Print this report

I forgot the unusual sensitivity spectrum for the high-frequency ranges on the night of August 13th, as shown in Figs. 1 and 2. They show a higher spectrum level than usual.

I don't know whether these events are related to this time FBL malfunction.

Images attached to this comment
shinji.miyoki - 18:41 Saturday 16 August 2025 (34848) Print this report

Photo 1 shows the master laser control panel when Miyakawa-kun noticed before touching the master laser and FBL amplifier. It clearly showed that the RED indicator, which means the master laser was off.

Just after taking this photo, photo 2 was taken and it showed 

  • seed: 1.814 (unit is 10mW)
  • temperature: 23.7C
  • current : 29.3 A

The RED indicator and the seed:1.8 are contradict with each other. However, as explained, the laser was actually emitted from the FBL at the power of 19W and its PMC refl image  also showed the existence of the laser beam.

Images attached to this comment
shinji.miyoki - 8:22 Sunday 17 August 2025 (34850) Print this report

We discussed the recovery plans on Sunday from 21:00 via zoom3.

  • Check PMC flashing with NeoLase. 
  • Check the master laser output power for the FBL.
  • IFO recovery using NeoLase might be better than FBL because of the following concerns.

The following are the concerned items

  • Is the present master laser for the FBL healthy? If not healthy, the same trouble can happen, even after replacing the FB amp part.
  • FBL should have one more interlock system to avoid FB amp operation without the seed light. It should be prepared in parallel with NeoLase operation.
shinji.miyoki - 10:41 Sunday 17 August 2025 (34854) Print this report

Spare lasers and amplifiers for IR

  • Amp (Nufern): Serial 8958 (completely the same model as the broken one. Mio-sensei verified its operation just before sending to KAGRA.)
    Model SUB-1174-2?
    Part No: PSFA-1064-10-40W-2-0
  • Amp (Nufern): Serial 80555 (bought 2018 Oct 11th, Weaker seed laser power version.)
    Product number: NUA-1064-PB-0050, 50W, PM amplifier, 1-15mW, collimator
    FDA #: 08R0003-014
    ECCN: EAR99
  • Laser: Mephisto 500FC for the Nufern FBL master
  • Laser: NeoLase in the PSL room: PRFPMI+RF lock was verified in the past.
shinji.miyoki - 12:32 Sunday 17 August 2025 (34855) Print this report

Miyakawa-kun tried the followings,

  • Measure the master laser output just before the Nufern FBL amplifier input.
    • ~ max 32 mW by using Newport Power meter (max 2W using NDO filter. Because he faced the power meter and fiber output by his hands, the measured power fluctuation was ~25mW ~ 32mW.
    • Timeline power measurement by using PDA10. No remarkable large fluctuation and oscillation.
  • Neolase operation
    • He found water leak in the piping route in the water chiller system because of the "known" reason. He will try to get a reliable part that is used in Toyama U today, if he can enter TU. Then he will retry on Monday.
Search Help
×

Warning

×